OpenStudy (anonymous):

Variable y varies directly with variable x, and y = 5 when x = 8. What is the value of x when y = 15?

OpenStudy (anonymous):

@Science_ALLY

OpenStudy (anonymous):

can you tag anyone thats is good in math????

OpenStudy (science_ally):

tag bibby

OpenStudy (anonymous):

@bibby

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

Same as the last prob i did with you y = k*x plug in x and y and solve for k

OpenStudy (godlovesme):

y directly varies with x means y=kx 5=k8 solve for k @pielover123

OpenStudy (anonymous):

x=8 y=5 k=45???

OpenStudy (danjs):

\[y = k*x~~~~5 = k*8~~~~~k = 5/8\]

OpenStudy (godlovesme):

 @pielover123 got it ?

OpenStudy (anonymous):

yes thanks:)

OpenStudy (danjs):

\[y = \frac{ 5 }{ 8 }*x\] put in y=15, solve for the X

OpenStudy (anonymous):

x=9

OpenStudy (anonymous):

right @Godlovesme

OpenStudy (godlovesme):



OpenStudy (anonymous):

24x thanks so much

OpenStudy (godlovesme):

welcome :)